# Machine-learning Guided Search for Phonon-mediated Superconductivity in Boron and Carbon Compounds

### v2

@@ I Introduction
Machine learning (ML) and Artificial intelligence (AI) are increasingly taking important roles for predicting materials properties including SC. Previous studies performing ML predictions based on random forest [49], regression [50, 51, 52, 53], classification [50], natural language processing (NLP) [54], and deep learning [55] models have trained on experimental data, mostly from SuperCon database [36]. Comparing to the more expensive and time consuming experiments to explore many new compounds to generate SC data for ML, the high throughput (HT) ab-initio first-principles approaches are valuable tools to obtain the data that can be trained to predict potential SC compounds. Several recent studies have utilized ab-initio computed data for training ML model and predicting SC. One approach involves performing BCS-inspired screening of materials to identify potential candidates based on certain key properties such as the Debye temperature and the density of states at the Fermi level (N(EF)) [9]. A study similar to that described in Ref. [9] has been conducted on a vast range of materials, but restricting the size of the compounds to eight or fewer atoms, as reported in Ref. [58]. Similarly, the ML model was developed to train and predict Eliashberg spectral functions using dynamically stable compounds [59]. In a recent study [60], a ML approach was used to predict the maximum Tc and corresponding pressure of binary metal hydrides. The input layer consisted of atomic properties of the heavier metallic atom, while the output layer had two nodes representing Tc and pressure. The ab-initio data utilized in this study has been collected from literature. Recently, ML-driven search with experimental feedback was also performed to discover a novel superconductor in Zr-In-Ni systems [61].
@@
Notably in these previous HT and ML studies, compounds of dynamical instability with imaginary phonon modes have been largely discarded. However, as shown by our recent EPC study [62] on Y2C3 with experimentally known Tc=18K[32], imaginary phonon of C dimer wobbly motion once stabilized can carry significant EPC contributions, which explains well the observed sizable Tc. In recent model Hamiltonian studies, phonon softening and anharmonicity have also been found to enhance Tc [63, 64]. Here we present a workflow that iteratively combines ab-initio calculations with an ML-guided search across the dataset of compounds with both dynamical stability and instability from imaginary phonon modes by focusing on boron/carbon/borocarbide (B/C/B+C) compounds. Ab-initio calculations were performed to compute the EPC strength (), the logarithmic average phonon frequency () and Tc of 417 compounds employing DFPT and isotropic Eliashberg approximation. Two major issues arise during DFPT calculations: choosing appropriate BZ sampling (k and q-mesh) for convergence and the problem of calculated dynamic instability. To address the convergence problem, we developed an ansatz test to check the convergence of EPC properties, particularly the Tc. For dynamically unstable compounds, we employed large electronic smearing, lattice distortion, and pressure to stabilize them. We then calculated their EPC properties, which were included in building the ML models. We evaluated ML models, specifically the crystal graph convolutional neural network (CGCNN) and the atomistic line graph neural network (ALIGNN), trained utilizing ab initio computed data to predict SC properties. Among the two models, especially when including the dynamically unstable compounds, ALIGNN consistently outperforms the CGCNN in predicting EPC properties. We predict a few promising SC compounds with formation energy just above the ground state convex hull. For dynamically stable systems, we predict TaNbC2 (28.4 K), Nb3B3C (16.4 K), Y2B3C2 (4.0 K) among others. For systems with dynamical instability and imaginary phonon modes, we predict Ca5B3N6 with a Tc as high as 35-42.4 K, besides Pd3CaB (7.0 K), and a few Ru compounds of MoRuB2 (15.6 K), RuVB2 (15.0 K), and RuSc3C4 (6.6 K).
@@ II.2 Overview of DFPT calculations
After the crystal structures of B and C compounds have been collected, the next step as shown in Fig.1 is to do HT calculations with DFPT on EPC properties for compounds with both known and unknown Tc using our recently developed high-throughput electronic structure package (HTESP) [157]. We performed DFPT calculations and computed EPC properties using the isotropic Eliashberg approximation. The accuracy of the EPC data is crucial for building reliable ML models. We have encountered two major challenges in the HT calculations with DFPT. One is the convergence with respect to BZ sampling and the other is dynamic instability. The first obstacle involved determining appropriate BZ samplings (k- and q- meshes) to compute the EPC properties, as these calculations become computationally expensive with dense meshes. To reduce the computational cost, we initiated an efficient screening by using a k-point mesh that accurately describes the ground-state structures and energetics. EPC quantities are then interpolated to fine k-mesh only twice the size of the coarse k-mesh. But we noticed that calculations with such grid combination can lead to inaccurate predictions, with discrepancies as high as approximately 10% of the total compounds with known Tc, giving NSC for known SC and vice versa. To address this problem, we developed an ansatz test to assess the convergence of Tc with respect to the k-point mesh. This approach leverages the decaying behavior of Tc with respect to Gaussian broadening width (), which is used in the double-delta integration. Initially, we acquired results using the DFPT method with the k-point mesh size from the MP database. Subsequently, we assessed the convergence of these results based on the convergence ansatz test. To ensure convergence, we repeated calculations with denser k-mesh for the cases where results did not pass this test. We applied this technique to the 113 dynamically stable compounds, and obtained reasonable accuracy for the calculated Tc with a mean absolute error (MAE) of 2.21 K compared to experimental data. Further details regarding the convergence with respect to the k- and q-point meshes are discussed in the Method section and SM. In addition to the 113 dynamically stable compounds with known Tc, we also computed the EPC properties of 268 compounds of unknown Tc with dynamical stability, resulting from the ML-guided search as summarized in Fig.1.
@@ II.3 Training and testing ML models
As shown in Fig. 1, we trained the initial ML models in Run 1 using a dataset of 250 dynamically stable compounds with 109 SC (calculated Tc 1 K) and 141 NSC (calculated Tc 1 K) including those 113 from SuperCon database that were already measured in experiments. This dataset encompasses 45 distinct SGs. For the purpose of evaluation, a separate collection of 58 stable compounds (18 SC and 40 NSC), belonging to 27 unique SGs (6 of which are not part of the training 45), was reserved exclusively for independent testing and was not involved in the ML training process. Moving from Run1 to Run2, we use the ML model from Run1 to predict Tc for the remaining dataset of B and C compounds. We then sort the predictions and select the top candidates of both SC and NSC for more DFPT calculations to close the ML-guided loop. In Run2, the original 250 stable systems were augmented with an additional 73 dynamically stable systems to refine the ML models. In the concluding stage of Run 3, we incorporated the results derived from the compounds with dynamic instability. Notably, this stage included results from 2 new SGs. Among the 36 results in this category, we incorporated 28 into the training set and added 8 compounds into the independent test set. In the overall count of 417 compounds with converged EPC properties, 181 were classified as SCs, whereas 236 were categorized as NSCs. The progress from Run 1 to Run 3 constitutes a loop, as illustrated in Fig. 1. In summary, our methodology has a series of iterations involving training and testing ML models with increasingly comprehensive datasets from dynamically stable to unstable compounds.
@@ II.4 Comparison of ML models
Here, It is important to note that the data first used do not include dynamically unstable cases. For Run 1, we trained ML models using 250 compounds, divided into training, validation, and testing sets in a ratio of 0.8:0.1:0.1. The mean absolute errors (MAEs) for the 10% test set in each training iteration are documented in SM Table-S6. An additional independent test set of 58 compounds was used to assess the predictability of the ML models and their MAE and predictions are plotted in Fig. 3. Apart from MAE, we also computed the R2 value for the ML parity plot, which is presented in Supplementary Table ST7. As expected, accurately predicting the exact EPC properties is challenging. However, the performance of the ML models improves when superconductors with stabilized phonons are included. The training process consisted of 3000 epochs using default settings provided by the ML packages, which have been thoroughly investigated in the original work [158, 159]. In both CGCNN and ALIGNN, training and validation procedures are employed. Checkpoints are established at regular intervals to store crucial parameters such as model weights and architecture. The ML packages operate automatically to retain and update the model exhibiting the best performance, determined by the lowest validation error. This iterative process also acts to mitigate overfitting concerns. Our Supporting Figure S12 and Supporting Table ST7 demonstrate that training up to 1500 epochs is sufficient to achieve the lowest validation error for both ALIGNN and CGCNN model. The performance of the models was evaluated by computing the MAE between the predicted and target quantities for the independent test set. The MAEs for CGCNN-predicted and stand at 0.23 and 105 K, respectively, while the corresponding numbers for ALIGNN are slightly higher at 0.28 and 113 K (Figs. 3(a) and (b)). In terms of predicting Tc, the CGCNN and ALIGNN models yield MAEs of 2.4 K and 3.8 K, respectively. Despite the relatively small magnitude of MAEs, the ML outcomes exhibit distinct clustering patterns, as shown in Fig. 3(c). Specifically, for the CGCNN model, the results tend to cluster closely along the “DFPT” axis (red arrow in Fig. 3(c)), whereas for ALIGNN, the clustering is pronounced along the “ML prediction” axis (blue arrow in Fig. 3(c)). An alternative approach, rather than directly training and predicting Tc, is to utilize ML-predicted values, specifically and , to estimate T using Eq. 1. This modification not only enhances predictive performance for both models but also slightly ameliorates the issue of clustering [Fig. 3(d)]. A comparable approach has been employed in a recent study [160], wherein and can be directly acquired from first principles calculations.
@@
As mentioned above, from the ML-guided search, we find quite some number of compounds with imaginary phonon and dynamical instability. In all the previous high throughput phonon-mediated SC studies, these dynamically unstable compounds are simply discarded, but as shown by our recent study [62] on Y2C3, some imaginary phonon modes once stabilized can carry a large EPC and give rise to a large Tc. Thus, specifically here we also include dynamically unstable compounds in ML. As far as we know, this is the first ML trained with both dynamically stable and unstable compounds for phonon-mediated SC. We carried out a distinct training and testing phase for ML models, incorporating dynamically unstable cases after stabilization, denoted as Run 3. These results predominantly included nonzero Tc. Among these, 28 results were added to the training dataset, while the remaining 8 were added for independent testing. In total, the independent test set now consists of the original 58 dynamically stable and additional 8 stabilized compounds with imaginary phonon modes. Both the training and test datasets were balanced across various SGs. The outcomes of Run 3 are illustrated in Fig. 4(a)-(d). In Run 3, the ALIGNN model consistently outperforms the CGCNN model across different superconducting properties. The ALIGNN model achieves MAEs of 0.27 for , 86 K for , 4.3 K for Tc, and 3.5 K for T. In contrast, the CGCNN model records MAEs of 0.34, 104 K, 4.5 K, and 5.3 K for the respective properties. The reason is that ALIGNN includes the bonding angles as part of training parameters, which better describe the dynamically unstable compounds, because imaginary phonon modes often involves lower-energy bond rotation with changing angle than higher-energy bond stretching vibration. Based on the information presented in Figs. 3 and 4, it is apparent that the ML model performs better in learning the parameters than and then Tc. As a result, it is justifiable to utilize Eq. 1 to estimate the ML-predicted critical temperature T rather than relying solely on the directly predicted Tc values. Furthermore, we separately computed the MAE for SC and NSC, as shown in Supplementary Table ST8. ALIGNN demonstrates better accuracy in predicting compared to CGCNN for both categories. However, it predicts more accurately than CGCNN for NSC, while performing slightly worse for SC. Additionally, a classification model was employed, with the results presented in the Supplementary Materials.
@@ II.5 Dynamically stable systems
In Table 1, we present the results of our ML-guided search for dynamically stable compounds with DFPT-calculated Tc exceeding 10 K, while the complete list of EPC properties of the 381 dynamically stable compounds are presented in SM from Table S7 to S13. It is worth noting that the compounds with high Tc tend to be thermodynamically metastable, as indicated by their formation energy significantly above the convex hull. B2CN in different phases, with up to 0.35 eV/atom, has been predicted to exhibit SC, consistent with earlier theoretical findings [161]. TaC and NbC, both have sizable Tc in the metastable hexagonal structure, while their more stable cubic structures in have already been observed with SC [162, 163]. In Table 1, the first compound close to the ground state convex hull ( 0.02 eV/atom) is TaNbC2, whose crystal structure and EPC properties are plotted in Fig. 5(a)-(c). It crystallizes in the trigonal structure of . The calculated EPC properties are , K, and K, with the majority of the contribution coming from phonons within the 3-6 THz range (Fig. 5(b) and (c)), as well as a significant contribution from the 16-20 THz range of C-dominated modes. Other compounds close to the ground state convex hull with 0.02 is TaB2. The presence of SC in TaB2 remains a subject of debate as discussed in the previous studies [95].

### v3

@@ Abstract
∗Corresponding author: Niraj K. Nepal, Email: nepalneeraz@gmail.com
†llw@ameslab.gov
@@ I Introduction
Machine learning (ML) and Artificial intelligence (AI) are increasingly taking important roles for predicting materials properties including SC. ML models usually offer a computationally inexpensive alternative to DFPT calculations for evaluating superconducting properties, completing predictions in minutes compared to the months, typically required for high throughput DFPT calculations of hundreds of compounds. Previous studies performing ML predictions based on random forest [49], regression [50, 51, 52, 53], classification [50], natural language processing (NLP) [54], and deep learning [55] models have trained on experimental data, mostly from SuperCon database [36]. Comparing to the more expensive and time consuming experiments to explore many new compounds to generate SC data for ML, the high throughput (HT) ab-initio first-principles approaches are valuable tools to obtain the data that can be trained to predict potential SC compounds. Several recent studies have utilized ab-initio computed data for training ML model and predicting SC. One approach involves performing BCS-inspired screening of materials to identify potential candidates based on certain key properties such as the Debye temperature and the density of states at the Fermi level (N(EF)) [9]. A study similar to that described in Ref. [9] has been conducted on a vast range of materials, but restricting the size of the compounds to eight or fewer atoms, as reported in Ref. [58]. More recently, a new ML model suitable for the limited size of EPC dataset has been developed to predict Eliashberg spectral functions rather than the overall EPC strength, which represents an important step to account for the correlation between phonon density of state and EPC spectral function [59]. In a recent study [60], a ML approach was used to predict the maximum Tc and corresponding pressure of binary metal hydrides. The input layer consisted of atomic properties of the heavier metallic atom, while the output layer had two nodes representing Tc and pressure. The ab-initio data utilized in this study has been collected from literature. Recently, ML-driven search with experimental feedback was also performed to discover a novel superconductor in Zr-In-Ni systems [61].
@@
Notably in these previous HT and ML studies, compounds of dynamical instability with imaginary phonon modes have been largely discarded. However, as shown by our recent EPC study [62] on Y2C3 with experimentally known Tc=18K[32], imaginary phonon of C dimer wobbly motion once stabilized can carry significant EPC contributions, which explains well the observed sizable Tc. In recent model Hamiltonian studies, phonon softening and anharmonicity have also been found to enhance Tc [63, 64]. Here we present a workflow that iteratively combines ab-initio calculations with an ML-guided search across the dataset of compounds with both dynamical stability and instability from imaginary phonon modes by focusing on boron/carbon/borocarbide (B/C/B+C) compounds. Ab-initio calculations were performed to compute the EPC strength (), the logarithmic average phonon frequency () and Tc of 417 compounds employing DFPT and isotropic Eliashberg approximation. Two major issues arise during DFPT calculations: choosing appropriate BZ sampling (k and q-mesh) for convergence and the problem of calculated dynamic instability. To address the convergence problem, we developed an ansatz test to check the convergence of EPC properties, particularly the Tc. For dynamically unstable compounds, we employed large electronic smearing, lattice distortion, and pressure to stabilize them. We then calculated their EPC properties, which were included in building the ML models. We evaluated ML models, specifically the crystal graph convolutional neural network (CGCNN) and the atomistic line graph neural network (ALIGNN), trained utilizing ab initio computed data to predict SC properties. The key new findings of this study are two-fold. First is the inclusion of the previously overlooked dynamical unstable compounds in the EPC training dataset, accounting for a sizable 20% of all compounds being studied, which would otherwise be omitted. Second, by including compounds with imaginary phonon modes, we reveal a clear performance gap between the two ML architectures, reinforcing the widely accepted principle that the completeness of a training dataset is as crucial as the ML model itself. Among the two models, especially when including the dynamically unstable compounds, ALIGNN consistently outperforms the CGCNN in predicting EPC properties. We predict a few promising SC compounds with formation energy just above the ground state convex hull. For dynamically stable systems, we predict TaNbC2 (28.4 K), Nb3B3C (16.4 K), Y2B3C2 (4.0 K) among others. For systems with dynamical instability and imaginary phonon modes, we predict Ca5B3N6 with a Tc as high as 35-42.4 K, besides Pd3CaB (7.0 K), and a few Ru compounds of MoRuB2 (15.6 K), RuVB2 (15.0 K), and RuSc3C4 (6.6 K).
@@ II.2 Overview of DFPT calculations
After the crystal structures of B and C compounds have been collected, the next step as shown in Fig.1 is to do HT calculations with DFPT on EPC properties for compounds with both known and unknown Tc using our recently developed high-throughput electronic structure package (HTESP) [157]. We performed DFPT calculations and computed EPC properties using the isotropic Eliashberg approximation. The accuracy of the EPC data is crucial for building reliable ML models. We have encountered two major challenges in the HT calculations with DFPT. One is the convergence with respect to BZ sampling and the other is dynamic instability. The first obstacle involved determining appropriate BZ samplings (k- and q- meshes) to compute the EPC properties, as these calculations become computationally expensive with dense meshes. To reduce the computational cost, we initiated an efficient screening by using a k-point mesh that accurately describes the ground-state structures and energetics. EPC quantities are then interpolated to fine k-mesh only twice the size of the coarse k-mesh. But we noticed that calculations with such grid combination can lead to inaccurate predictions, with discrepancies as high as approximately 10% of the total compounds with known Tc, giving NSC for known SC and vice versa. To address this problem, we developed an ansatz test to assess the convergence of Tc with respect to the k-point mesh. This approach leverages the decaying behavior of Tc with respect to Gaussian broadening width (), which is used in the double-delta integration. Initially, we acquired results using the DFPT method with the k-point mesh size from the MP database. Subsequently, we assessed the convergence of these results based on the convergence ansatz test. To ensure convergence, we repeated calculations with denser k-mesh for the cases where results did not pass this test. We applied this technique to the 113 dynamically stable compounds, and obtained reasonable accuracy for the calculated Tc with a mean absolute error (MAE) of 2.21 K compared to experimental data, as presented in Supplementary Figure S2 (b). Moreover, incorporating dynamically unstable compounds further improves the dataset quality, as evidenced in supplementary Fig. S2(c)–(d) by the enhanced score in the comparison with experiment. Further details regarding the convergence with respect to the k- and q-point meshes are discussed in the Method section and SM. In addition to the 113 dynamically stable compounds with known Tc, we also computed the EPC properties of 268 compounds of unknown Tc with dynamical stability, resulting from the ML-guided search as summarized in Fig.1.
@@ II.3 Training and testing ML models
As shown in Fig. 1, we trained the initial ML models in Run 1 using a dataset of 250 dynamically stable compounds with 109 SC (calculated Tc 1 K) and 141 NSC (calculated Tc 1 K) including those 113 from SuperCon database that were already measured in experiments. This dataset encompasses 45 distinct SGs. For the purpose of evaluation, a separate collection of 58 stable compounds (18 SC and 40 NSC), belonging to 27 unique SGs (6 of which are not part of the training 45), was reserved exclusively for independent testing and was not involved in the ML training process. Moving from Run 1 to Run 2, we use the Run 1 ML model (trained on 250 dynamically stable compounds) to predict Tc for the remaining B–C compounds, which is fewer than 450, because some compounds were already found to be dynamically unstable but without EPC data during Run 1 ML training. We then sort the predictions and select the top candidates of both SC and NSC for more DFPT calculations to close the ML-guided loop. In Run2, the original 250 stable systems were augmented with an additional 73 dynamically stable systems to refine the ML models. In the concluding stage of Run 3, we incorporated the results derived from the compounds with dynamic instability. Notably, this stage included results from 2 new SGs. Among the 36 results in this category, we incorporated 28 into the training set and added 8 compounds into the independent test set. In the overall count of 417 compounds with converged EPC properties, 181 were classified as SCs, whereas 236 were categorized as NSCs. The progress from Run 1 to Run 3 constitutes a loop, as illustrated in Fig. 1. In summary, our methodology has a series of iterations involving training and testing ML models with increasingly comprehensive datasets from dynamically stable to unstable compounds.
@@ II.4 Comparison of ML models
Here, It is important to note that the data first used do not include dynamically unstable cases. For Run 1, we trained ML models using 250 compounds, divided into training, validation, and testing sets in a ratio of 0.8:0.1:0.1. The mean absolute errors (MAEs) for the 10% test set in each training iteration are documented in Supplementary Table S1. An additional independent test set of 58 compounds was used to assess the predictability of the ML models and their MAE and predictions are plotted in Fig. 3. Apart from MAE, we also computed the R2-score for the ML parity plot, which is presented in Supplementary Table S2. As expected, accurately predicting the exact EPC properties is challenging. However, the performance of the ML models improves when superconductors with stabilized phonons are included. The training process consisted of 3000 epochs using default settings provided by the ML packages, which have been thoroughly investigated in the original work [158, 159]. In both CGCNN and ALIGNN, training and validation procedures are employed. Checkpoints are established at regular intervals to store crucial parameters such as model weights and architecture. The ML packages operate automatically to retain and update the model exhibiting the best performance, determined by the lowest validation error. This iterative process also acts to mitigate overfitting concerns. Our Supporting Figure S3 and Supporting Table S2 demonstrate that training up to 1500 epochs is sufficient to achieve the lowest validation error for both ALIGNN and CGCNN model. The performance of the models was evaluated by computing the MAE between the predicted and target quantities for the independent test set. The MAEs for CGCNN-predicted and stand at 0.23 and 105 K, respectively, while the corresponding numbers for ALIGNN are slightly higher at 0.28 and 113 K (Figs. 3(a) and (b)). In terms of predicting Tc, the CGCNN and ALIGNN models yield MAEs of 2.4 K and 3.8 K, respectively. Despite the relatively small magnitude of MAEs, the ML outcomes exhibit distinct clustering patterns, as shown in Fig. 3(c). Specifically, for the CGCNN model, the results tend to cluster closely along the “DFPT” axis (red arrow in Fig. 3(c)), whereas for ALIGNN, the clustering is pronounced along the “ML prediction” axis (blue arrow in Fig. 3(c)). An alternative approach, rather than directly training and predicting Tc, is to utilize ML-predicted values, specifically and , to estimate T using Eq. 1. This modification not only enhances predictive performance for both models but also slightly ameliorates the issue of clustering [Fig. 3(d)]. A comparable approach has been employed in a recent study [160], wherein and can be directly acquired from first principles calculations.
@@
As mentioned above, from the ML-guided search, we find quite some number of compounds with imaginary phonon and dynamical instability. In all the previous high throughput phonon-mediated SC studies, these dynamically unstable compounds are simply discarded, but as shown by our recent study [62] on Y2C3, some imaginary phonon modes once stabilized can carry a large EPC and give rise to a large Tc. Thus, specifically here we also include dynamically unstable compounds in ML. As far as we know, this is the first ML trained with both dynamically stable and unstable compounds for phonon-mediated SC. We carried out a distinct training and testing phase for ML models, incorporating dynamically unstable cases after stabilization, denoted as Run 3. These results predominantly included nonzero Tc. Among these, 28 results were added to the training dataset, while the remaining 8 were added for independent testing. In total, the independent test set now consists of the original 58 dynamically stable and additional 8 stabilized compounds with imaginary phonon modes. Both the training and test datasets were balanced across various SGs. The outcomes of Run 3 are illustrated in Fig. 4(a)-(d). In Run 3, the ALIGNN model consistently outperforms the CGCNN model across different superconducting properties. The ALIGNN model achieves MAEs of 0.27 for , 86 K for , 4.3 K for Tc, and 3.5 K for T. In contrast, the CGCNN model records MAEs of 0.34, 104 K, 4.5 K, and 5.3 K for the respective properties. The reason is that ALIGNN includes the bonding angles as part of training parameters, which better describe the dynamically unstable compounds, because imaginary phonon modes often involves lower-energy bond rotation with changing angle than higher-energy bond stretching vibration. Based on the information presented in Figs. 3 and 4, it is apparent that the ML model performs better in learning the parameters than and then Tc. As a result, it is justifiable to utilize Eq. 1 to estimate the ML-predicted critical temperature T rather than relying solely on the directly predicted Tc values. Furthermore, we separately computed the MAE for SC and NSC, as shown in Supplementary Table S3. ALIGNN demonstrates better accuracy in predicting compared to CGCNN for both categories. However, it predicts more accurately than CGCNN for NSC, while performing slightly worse for SC. Additionally, a classification model was employed, with the results presented in the Supplementary Table S4.
@@ II.5 Dynamically stable systems
In Table 1, we present the results of our ML-guided search for dynamically stable compounds with DFPT-calculated Tc exceeding 10 K, while the complete list of EPC properties of the 381 dynamically stable compounds are presented in SM from Table S10 to S17. It is worth noting that the compounds with high Tc tend to be thermodynamically metastable, as indicated by their formation energy significantly above the convex hull. B2CN in different phases, with up to 0.35 eV/atom, has been predicted to exhibit SC, consistent with earlier theoretical findings [161]. TaC and NbC, both have sizable Tc in the metastable hexagonal structure, while their more stable cubic structures in have already been observed with SC [162, 163]. In Table 1, the first compound close to the ground state convex hull ( 0.02 eV/atom) is TaNbC2, whose crystal structure and EPC properties are plotted in Fig. 5(a)-(c). It crystallizes in the trigonal structure of . The calculated EPC properties are , K, and K, with the majority of the contribution coming from phonons within the 3-6 THz range (Fig. 5(b) and (c)), as well as a significant contribution from the 16-20 THz range of C-dominated modes. Other compounds close to the ground state convex hull with 0.02 is TaB2. The presence of SC in TaB2 remains a subject of debate as discussed in the previous studies [95].